# Labelling Data with Unknown References

### v3

@@ Abstract
An evaluator is trustworthy when there exists some agreed-upon way to measure its performance as a labeller. The two ways to establish trustworthiness are either by testing it, or by assuming the evaluator ‘knows’ somehow the way to label the corpus. However, if labelled references (e.g., a development set) are unavailable, neither of these approaches work: the former requires the data, and the latter is an assumption, not evidence. To address this, we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without any existing references. Our algorithm works by successively posing challenges to said evaluator. We show that this is sufficient to establish trustworthiness w.h.p., in such a way that when the evaluator actually knows the way to label the corpus, the No-Data Algorithm accepts its output; and, conversely, flags untrustworthy evaluators when these are unable to prove it. We present formal proofs of correctness, empirical tests, and applications to LLMs-as-judges on low-resource languages.
@@ 1 Introduction
One of the most fundamental problems in AI is the question of measurement. It has been known for a while that some metrics have a certain level of unreliability; for example, by not corresponding to human judgements appropriately (Reiter, 2018; Gehrmann et al., 2023; Bavaresco et al., 2024; Hada et al., 2024). With the rise of LLMs as evaluators, typically known as LLMs-as-judges, scaling measurement to multiple subproblems has become ubiquitous (de Wynter, 2025). However, their reliability and trustworthiness as a measurement tool has led to a substantial divide in the research community, and specifically around the validity of some results (Gehrmann et al., 2023; Rabinovich and Anaby Tavor, 2025; Moghe et al., 2023).
@@
A key component of the question of measurement, and the focus of our work, is whether the tool providing the evaluation (the evaluator) may be trusted. This trust is typically established with some sort of ground truth, such as an existing set of labelled references with which to perform statistical analyses (e.g., a $t$-test), or a benchmark with previously-published results. Due to uniqueness of the problem, cost of annotation, or other reasons, references may not readily exist, or be too scarce to be statistically significant. Likewise, contamination of public benchmarks has become a major concern (see, e.g., Sainz et al. 2023). It follows then that the only way one can scientifically establish trust on an evaluator whose capabilities are unknown, in a no-reference scenario, is with a formal proof.
In this paper we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without previously-labelled data. The algorithm is designed to work in-line with contemporary machine learning techniques–namely, LLMs-as-judges–where an evaluator claims it knows the map between datapoints and labels. It is up to a verifier to decide, based on the outcome of certain challenges, whether the evaluator may be trusted. The No-Data Algorithm requires a linear number of calls and does not require labelled datasets (e.g., development or training sets) or knowledge of the labels in the test set to establish trust.
We show that our algorithm, after $r$ calls, bounds confidence on the evaluator to a $(1/4)^{r}$ probability of error. This success rate reflects the evaluator’s knowledge–or lack thereof: if it knows how to label the set, the accuracy and success rate, as reported by the No-Data Algorithm, will be high. Conversely, if the evaluator lies about knowing the labels, the accuracy could be high–perhaps by luck–but the success rate will not. We present empirical tests with traditional machine-learning (i.e., a decision tree) and contemporary (r. LLMs) evaluation methods. We also showcase an application of the No-Data Algorithm by judging an LLM-as-a-judge’s ability to label a dataset in West Frisian, a low-resource language. Both experiments adjust to the theory’s predictions. [^1] Code and data may be found in https://github.com/adewynter/no_data_algorithm. Our results demonstrate that trust in an evaluator may be established via formal methods without the need for references, and enables their application in scenarios where labels may be extremely scarce, such as market research, low-resource languages (the subject of our experiments), and the medical domain.
@@ 2 Related Work
More generally, it is known that metric choice, and hence the choice of evaluator, is system-dependent (Chen et al., 2024; Moghe et al., 2023; Flamich et al., 2025; von Däniken et al., 2024; Pombal et al., 2025). Poorly-developed rubrics also lead to unreliable human judgements (Clark et al., 2021; van der Lee et al., 2019), and thus the evaluator trustworthiness problem is also a rubric problem. There are algorithms (Northcutt et al., 2017; Zhang and Sabuncu, 2018) and paradigms (Northcutt et al., 2021) for learning with noisy teachers. Likewise, provable methods have been used with success in other areas, such as representation learning (Jovanović et al., 2023), and to improve robustness via synthetic data generation (Dimitrov et al., 2022; Fischer et al., 2022). However, they all assume that a (perhaps noisily) labelled dataset exists.
Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans (Open AI, 2023; Chiang et al., 2024; Chiang and Lee, 2023; Liu et al., 2023; Wei et al., 2024; Zheng et al., 2023). Some of these criticisms are that they do not correlate with human judgements well (Bavaresco et al., 2024; Hada et al., 2024; Chen et al., 2024; de Wynter et al., 2025; Tjuatja et al., 2024); are sensitive to their prompt (Lu et al., 2022; Hida et al., 2024; Ye and Durrett, 2022); their performance could be the product of memorisation (Lee et al., 2023; de Wynter et al., 2023; Sainz et al., 2023); and even that they do not understand the task at all (de Wynter and Yuan, 2024; Webson and Pavlick, 2022; Ye and Durrett, 2022). Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics (Schaeffer et al., 2023; Chen et al., 2024), to findings that their output reasoning steps contain spurious reasoning (Turpin et al., 2023; Lanham et al., 2023; Wu et al., 2024). Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs (Li et al., 2024).
The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol is based off the well-known Arthur-Merlin (AM) protocol from Babai (1985). It is a type of zero-knowledge proof; see Goldreich (2009) and Arora and Barak (2009) for primers on the subject. That said, the EV protocol is different in many regards to the AM protocol, since it is designed to fit arbitrary inputs and the larger No-Data Algorithm. In particular, the EV protocol always reveals its secrets, which sets it apart from applications in theoretical computer science. Likewise, our use of the rubric and the aggregator functions are generalisations to partial functions that are not present (or needed) in previous works.
@@ 3.1 Notation
Our work relates to datapoints $x\in X$ and labels $y\in Y$. Whenever the label set $Y$ is unspecified, it is the binary label set $Y=\{0,1\}$. We also let $X\subset\{0,1\}^{n}$; that is, every datapoint is a $n$-bit binary string modelling a phenomenon. The assumptions on $X$ and $Y$ are to simplify our proofs; see Section 7 for a discussion on their relevancy. For a binary string $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$. The construction of this specific set will depend on the setup, and will be discussed in detail in Section 4.1.
@@ 3.2 Definitions
In line with standard learning theory (Valiant, 1984), we assume that there exists an unknown map $f\colon X\rightarrow Y$ that takes datapoints to labels faithfully–that is, $f$ provides the true labelling. Our task is to determine whether a function $E\colon X\rightarrow Y$ (a labeller, or evaluator) is equivalent to $f$, without any knowledge of the map itself. This is measured with its error rate, or, conversely, its accuracy.
We also introduce four definitions which we will use in our work:
A criterion $c\colon X\rightarrow\{0,1\}$ maps a datapoint $X$ to a single bit.
A rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ maps the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ to a bitstring that is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$. We say a rubric is total if it explicitly decomposes nonlinear criteria with arity >1>1 (e.g. the xor $c_{i}=c_{a}\oplus c_{b}$), and write it as $\bar{\mathcal{C}}$. When the evaluation of said criteria can be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
@@
Finally, for two $x,x^{\prime}\in X$, we say that $x$ is similar to $x^{\prime}$ if both $x\cong x^{\prime}$ and $x\equiv x^{\prime}$.
Intuitively, the rubric is the set of reasons why a datapoint $x$ would take a given label (for example, an annotation rubric passed in to human labellers, the last layer in a model, etc.); while the aggregator decides the final label based on that rubric (e.g., via an activation function, a majority vote, etc.). Natural-language examples are in Appendix E.
We are now ready to introduce the core assumption of our work:
Assumption 1.
@@
It is not a strong assumption: after all, datapoints must have a reason to acquire a label. When there is more than one reason, there must be a way to decide how to aggregate them. However, this decomposition is crucial to our work.
@@ 4 The No-Data Algorithm
The No-Data Algorithm is an algorithm with two components (players): the evaluator and the verifier.Throughout the algorithm, for every datapoint $x\in X$, the evaluator must convince the verifier that it is able to label $x$ correctly (i.e., that it knows $f$).
@@
The No-Data Algorithm then flips the label with probability $\phi$ in case of failure. After observing every $x\in X$, it returns the success count and final labelling. The evaluator’s trustworthiness (and hence the final labelling’s) is established with the success count.
The full algorithm is in Algorithm 1. We and provide a characterisation of the No-Data Algorithm (bounds, correctness, etc.) in Section 5.
@@ 4.1 The Evaluator-Verifier (EV) Protocol
In the EV protocol both players are given a datapoint $x\in X$, rubric $C$, and nothing else. In particular, $\sigma$ is not explicitly given to either player. The evaluator claims to know the map $f\colon X\rightarrow Y$. Since $C$ is given, the claim is then that it has prior knowledge of $\sigma$. The verifier’s goal is to be sufficiently sure that the evaluator’s choice of label $x$ for a given $x\in X$ may be trusted.
@@ 4.1.1 Setup
The EV protocol is played for $r$ rounds. At every round, the evaluator generates a similar datapoint $x^{\prime}$ to $x$, and a ‘partial’ label $\tilde{y}^{\prime}$ that–supposedly–is equal to $y$. It is partial because there is no access to $f$, and hence the best the evaluator can do is to guess $\tilde{y}^{\prime}$ based on $C(x^{\prime})$. Once it has generated the tuple $\langle x^{\prime},\tilde{y}^{\prime}\rangle$, the evaluator may not change it in this round. The verifier poses one of two challenges, selected uniformly at random. Either challenge measures equality between $x^{\prime}$ and $x$, either up to isomorphism or up to permutation. If the evaluator fails, the protocol terminates and returns failure. Otherwise, it starts again. Either in success or failure, the protocol also returns $\tilde{y}^{\prime}$. See Figure 1 for a diagram of the flow of the EV protocol.
The verifier challenges are central to the work, and perform both structural and valuation checks:
@@
$\forall s\in S_{x}\exists t\in S_{x^{\prime}}\,.\,\forall c\in\bar{\mathcal{C}}\,c(s)=c(t);$ (2)
@@
Remark 1.
Since the challenges are public, the evaluator could cheat and generate an $x^{\prime}$ to pass one of the challenges, but not both. Otherwise, the implication is that it knows $f$, and it will pass the challenges every time. This is why the challenges assert both structure ($x\equiv x^{\prime}$) and final encoding ($x\cong x^{\prime}$). This observation is central to our correctness proofs (Section 5).
5 Characterisation
Lemma 1 (EV Protocol Correctness Bound).
@@ 5 Characterisation
(Sketch; a complete proof is in Appendix A.1). We separate the proof in two parts: whether the challenges are necessary and sufficient to determine whether the evaluator is lying; and the probability bound to which it happens.
On the first part, we note that Challenge 1 determines the internal structure of the string, but does not account for the valuation of $C(x^{\prime})$. This challenge is necessary because two $x,x^{\prime}$ could lead to the same evaluation of some $c\in\mathcal{C}$ (e.g., disjunctive clauses), and that would not mean that they are the same up to permutation.
However, Challenge 2 requires strict equality in $C(x^{\prime})=C(x)$, and it is indeed order-dependent, but does not account for said internal structure. Hence both challenges provide some, but not all, information about $x,x^{\prime}$ and the claimed $f(x),f(x^{\prime})$. They are then both sufficient and necessary to determine if an evaluator is lying. More importantly, they are mutually exclusive in the information they provide. This is needed because a lying evaluator could know ahead of time the challenges.
For the second part we handle two cases where the evaluator could lie (or cheat): either (1) it can generate an appropriate $x^{\prime}$, but not a correct $C(x^{\prime})$; or (2) it provides a correct $C(x^{\prime})$ but an incorrect $x^{\prime}$. In either case, the probability that it gets lucky and/or passes the challenge by cheating is $1/4$; after $r$ rounds, it is $(1/4)^{r}$. ∎
Theorem 1 (No-Data Algorithm Correctness Bound).
@@
(Sketch; a complete proof is in Appendix A.4). The accuracy of any evaluator is directly tied to its success in the algorithm: if it knows the label, then the checks by the EV protocol will pass. Even under the constraint that the mislabelled datapoints are tied to the evaluator’s lies, errors could still get past undetected. From Lemma 1, we know the error bound for every label flip, and therefore the probability of mislabelling the entire dataset. We may compute the latter via an application of the union bound, that is, $\Pr[\text{error}]\leq(1-\alpha)(1-\phi+\phi(1/4)^{r})$. ∎
The implication of Theorem 1 is that a good evaluator with relatively high accuracy and success rate, and a well-calibrated $\phi$, will have a true accuracy higher than what the No-Data Algorithm shows. This means that reliable evaluators, as per the algorithm, may be used with the expectation that their ‘true’ performance will be better. Another interpretation of this result is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier.
It is worth mentioning that the assumption for Theorem 1 around always lying when mislabelling (and hence always being wrong) are not necessarily applicable to all evaluators, since a particularly tricky evaluator could still guess the right label and yet not pass the challenges. We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust, especially in scenarios without any references.

### v4

# Algorithmically Establishing Trust in Evaluators

@@ Abstract
An evaluator, such as an LLM-as-a-judge, is trustworthy when there exists some agreed-upon way to measure its performance as a labeller. Traditional approaches either rely on testing the evaluator against references or assume that it ‘knows’ somehow the correct labelling. Both approaches fail when references are unavailable: the former requires data, and the latter is an assumption, not evidence. To address this, we introduce the ‘No-Data Algorithm’, which provably establishes trust in an evaluator without requiring any labelled data. Our algorithm works by successively posing challenges to said evaluator. We prove that after $r$ challenge rounds, it accepts an evaluator which knows the correct labels with probability $\geq 1-(1/4)^{r}$, and reliably flags untrustworthy ones. We present formal proofs of correctness, empirical tests, and applications to assessing trust in LLMs-as-judges for low-resource language labelling. Our work enables scientifically-grounded evaluator trust in low-data domains, addressing a critical bottleneck for scalable, trustworthy LLM deployment.
evaluation, trust, LLMs-as-judges
@@ 1 Introduction
One of the most fundamental problems in AI is the question of measurement. Many commonly used metrics have a certain level of unreliability; for example, by not corresponding to human judgements (Reiter, 2018; Gehrmann et al., 2023; Bavaresco et al., 2024; Hada et al., 2024). With the rise of LLMs as evaluators, typically known as LLMs-as-judges, scaling measurement to multiple subproblems has become ubiquitous (De Wynter, 2025). However, their reliability and trustworthiness as measurement tools has led to division within the research community, specifically, about the validity of some results (Gehrmann et al., 2023; Rabinovich & Anaby Tavor, 2025; Moghe et al., 2023).
@@
A key issue on the question of measurement is whether the evaluation tool (the evaluator) may be trusted. This trust is typically established with a ground truth artifact, such as a labelled dataset for statistical analysis, or a benchmark. However, references may not exist or be scarce, due to annotation cost or uniqueness of the problem. In this case, statistical conclusions would not be meaningful. Even when references exist, contamination of benchmarks has become a major concern (see, e.g., Sainz et al. 2023). Hence, when no references exist, scientifically establishing trust on an evaluator can only be done with a formal proof.
We introduce an algorithm (the ‘No-Data Algorithm’) which allows trust assessment without any reference data. It is designed for contemporary settings, such as LLMs-as-judges, where an evaluator claims knowledge of the correct labels, and constrained to a binary classification setting. A verifier then decides, based on the evaluator’s successful responses to specific challenges, whether it is trustworthy. This requires a linear number of queries, no labelled datasets (e.g., development or training sets), or prior knowledge of the labels. The method is inspired by zero-knowledge proof protocols, such as two-factor authentication.
1.1 Contributions
Our contributions are theoretical and experimental. Theoretically, we prove that after $r$ calls, the No-Data Algorithm establishes trustworthiness on the evaluator to a $(1/4)^{r}$ probability of error. This measure robustly reflects whether the evaluator actually has label knowledge. A knowledgeable evaluator will have high accuracy and success rate. Conversely, a lying evaluator [^1] In this work we use ‘lying’ to denote both uninformed and sycophantic evaluators, and we evaluate both cases. may achieve high accuracy only by chance, but will not attain a high success rate.
Empirically, we test the No-Data Algorithm with traditional machine-learning (i.e., a decision tree) and contemporary (r. LLMs) evaluators. We demonstrate an application to assessing an LLM-as-a-judge’s trustworthiness in a novel dataset in West Frisian, a low-resource language. We also include ablation studies with LLMs-as-judges, prompts, algorithm components, and extend our approach to $k$-ary label sets. [^2] All data is in https://github.com/adewynter/no_data_algorithm.
Our work shows that trust in an evaluator may be established formally without references, enabling scientific evaluation in domains with extremely scarce labels, such as market research, low-resource languages (the subject of our experiments), medicine. To our knowledge, ours is the first mathematically rigorous approach to establishing trust in an LLM-as-a-judge without labelled references.
@@ 2 Related Work
More generally, it is known that metric choice, and hence the choice of evaluator, is system-dependent (Chen et al., 2024; Moghe et al., 2023; Flamich et al., 2025; von Däniken et al., 2024; Pombal et al., 2025). Poorly-developed rubrics also lead to unreliable human judgements (Clark et al., 2021; van der Lee et al., 2019), and thus the evaluator trustworthiness problem is also a rubric problem. There are algorithms (Northcutt et al., 2017; Zhang & Sabuncu, 2018) and paradigms (Northcutt et al., 2021) for learning with noisy teachers. Likewise, provable methods have been used with success in other areas, such as representation learning (Jovanović et al., 2023), and to improve robustness via synthetic data generation (Dimitrov et al., 2022; Fischer et al., 2022). However, they all assume that a (perhaps noisily) labelled dataset exists.
Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans (Open AI, 2023; Chiang et al., 2024; Chiang & Lee, 2023; Liu et al., 2023; Wei et al., 2024; Zheng et al., 2023). Some of these criticisms are that they do not correlate with human judgements well (Bavaresco et al., 2024; Hada et al., 2024; Chen et al., 2024; De Wynter et al., 2025; Tjuatja et al., 2024); are sensitive to their prompt (Lu et al., 2022; Hida et al., 2024; Ye & Durrett, 2022); their performance could be the product of memorisation (Lee et al., 2023; De Wynter et al., 2023; Sainz et al., 2023); and even that they do not understand the task at all (De Wynter & Yuan, 2024; Webson & Pavlick, 2022; Ye & Durrett, 2022). Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics (Schaeffer et al., 2023; Chen et al., 2024), to findings that their output reasoning steps contain spurious reasoning (Turpin et al., 2023; Lanham et al., 2023; Wu et al., 2024). Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs (Li et al., 2024).
The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol, is based off the well-known Arthur-Merlin (AM) protocol from Babai (1985). It is a type of zero-knowledge proof; see Goldreich (2009) and Arora & Barak (2009) for primers on the subject. That said, the EV protocol is different in many regards to the AM protocol, since it is designed to fit arbitrary inputs and the larger No-Data Algorithm. In particular, it does not provide privacy or secrecy as other cryptographic protocols do; instead, we adapt the challenge-response game to evaluation. Likewise, our use of the rubric and the aggregator functions are generalisations to partial functions that are not present (or needed) in previous works.
@@ 3.1 Notation
Our work relates to datapoints $x\in X$ and labels $y\in Y$. Whenever the label set $Y$ is unspecified, it is the binary label set $Y=\{0,1\}$. We also let $X\subset\{0,1\}^{n}$; that is, every datapoint is a $n$-bit binary string modelling a phenomenon. The assumptions on $X$ and $Y$ are to simplify our proofs; see Section 7 for a discussion on their relevancy. Experimentally, $X$ may be other (concrete) data types, such as natural language strings. For an $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$. The construction of this specific set will depend on the setup, and will be discussed in detail in Section 4.1.
@@ 3.2 Definitions
In line with standard learning theory (Valiant, 1984), we assume that there exists an unknown map $f\colon X\rightarrow Y$ taking datapoints to labels faithfully–that is, $f$ provides the true labelling. Our task is to determine whether a function $E\colon X\rightarrow Y$ (a labeller, or evaluator) is equivalent to $f$, without any knowledge of the map itself. This is measured with its error rate, or, conversely, its accuracy.
We also introduce the following definitions:
A criterion $c\colon X\rightarrow\{0,1\}$ classifies datapoints.
A rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ maps the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ to a bitstring which is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$. We say a rubric is total if it explicitly decomposes nonlinear criteria with arity >1>1 (e.g. the xor $c_{i}=c_{a}\oplus c_{b}$), and write it as $\bar{\mathcal{C}}$. When the evaluation of said criteria may be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
@@
For two $x,x^{\prime}\in X$, we say that $x$ is similar to $x^{\prime}$ if both $x\cong x^{\prime}$ and $x\equiv x^{\prime}$.
Finally, in line with the AM protocol, we use lying to refer to any evaluator that presents unsubstantiated claims of correctness, including both uninformed (random guessing) and sycophantic (overconfident without knowledge) behaviour. The difference is that the latter could perform well in-distribution, but fail otherwise. The former, on the other hand, will simply not know how to solve the problem and fail at both. Without labels it is difficult to discern both cases, although we show that our algorithm is capable of it.
Remark that, intuitively, the rubric is the set of reasons why a datapoint belongs to a certain label class (e.g, an annotation rubric passed in to human labellers). The label is thus based on evaluating the rubric and aggregating its criteria with $\sigma$ (e.g., with a majority vote). Natural-language examples are in Appendix E.1; experiment with potential rubric weaknesses in Section 6.2; and discuss them further in Section 7.
We are now ready to introduce our main assumption:
Assumption 3.1.
@@
This is not a strong assumption: after all, datapoints must have a reason (criterion) to acquire a label. This reason must be clearly stated; and, when there is more than one reason, there must be a way to decide how to aggregate them. In realistic scenarios, $C$ or $\sigma$ may not be known. There, our method may still be applied with user-defined surrogates, but derived trust bounds are with respect to these surrogates rather than any ‘true’ but unknown ground truth. We come back to this in Section 7.
@@ 4 The No-Data Algorithm
The No-Data Algorithm is an algorithm with two components, or players: the evaluator and the verifier. Throughout the algorithm’s run, for every datapoint $x\in X$, the evaluator must convince the verifier that its choice of label for $x$ is correct (i.e., that it knows $f$).
@@
The No-Data Algorithm then flips the label with probability $\phi$ in case of failure. After observing all $X$, it returns the success count and final labelling. The evaluator’s trustworthiness (and hence the final labelling’s) is established via the average success rate.
The full algorithm is in Algorithm 1. We provide proofs of correctness in Section 5.
@@ 4.1 The Evaluator-Verifier (EV) Protocol
In the EV protocol both players are given a datapoint $x\in X$ and rubric $C$. The aggregator $\sigma$ is not given. The evaluator claims to know $f\colon X\rightarrow Y$; and, since $C$ is known, the claim reduces to the evaluator having knowledge of $\sigma$. The verifier’s goal is thus to be sufficiently sure that the evaluator’s labelling of a given $x\in X$ may be trusted.
@@ 4.1.1 Setup
The EV protocol is played for $r$ rounds. At every round, the evaluator generates a similar datapoint $x^{\prime}$ to $x$, and a ‘partial’ label $\tilde{y}^{\prime}$ that–supposedly–is equal to yy. [^3] Remark that the process by which the evaluator comes up with $x^{\prime}$ does not need to be generation in the LLM sense. It is instead left up to the evaluator, not unlike other zero-knowledge proofs. It is partial because there is no access to $f$, and hence the best the evaluator can do is to guess $\tilde{y}^{\prime}$ based on $C(x^{\prime})$. Once it has generated the tuple $\langle x^{\prime},\tilde{y}^{\prime}\rangle$, the evaluator may not change it in this round. The verifier poses one of two challenges, selected uniformly at random. Both challenges measure equality between $x^{\prime}$ and $x$, either up to isomorphism (validating the label) or up to permutation (structurally validating the criteria). If the evaluator fails, the protocol terminates and returns failure. Otherwise, it starts over. Regardless of the outcome, the protocol also returns $\tilde{y}^{\prime}$. See Figure 1 for a diagram of the protocol.
@@
$\forall s\in S_{x},\exists t\in S_{x^{\prime}}\,.\,\forall c\in\bar{\mathcal{C}},\,c(s)=c(t);$ (2)
@@
Remark 4.1.
Since the challenges are public, the evaluator could cheat and generate an $x^{\prime}$ to pass one of the challenges, but not both. Otherwise, the implication is that it knows $f$, and it will pass the challenges every time. This is why the challenges assert both structure ($x\equiv x^{\prime}$) and final encoding ($x\cong x^{\prime}$). This observation is central to our correctness proofs from Section 5.
5 Correctness
Lemma 5.1 (EV Protocol Correctness Bound).
@@ 5 Correctness
(Sketch; a complete proof is in Appendix A.1). We split the proof in two: (1) whether the challenges are necessary and sufficient to detect if the evaluator is lying; and (2) the probability bound to which this detection occurs.
For the first part, note that Challenge 1 determines the internal structure of the string, but does not account for the valuation of $C(x^{\prime})$. This challenge is necessary because two $x,x^{\prime}$ could lead to the same evaluation of some $c\in\mathcal{C}$ (e.g., disjunctive clauses), yet this does not imply that they are equal up to permutation.
However, Challenge 2 requires strict equality in $C(x^{\prime})=C(x)$. It is indeed order-dependent, but does not account for internal structure. Hence both challenges provide some, but not all, information about $x,x^{\prime}$ and the claimed $f(x),f(x^{\prime})$. Thus they are together sufficient and necessary to determine if an evaluator is lying.
Since they are mutually exclusive in the information they provide, it follows that an evaluator passing both challenges will not be lying. Thus, a lying evaluator may only be able to pass one challenge. Hence we handle two cases where the evaluator could lie (or cheat): either (1) it can generate an appropriate $x^{\prime}$, but not a correct $C(x^{\prime})$; or (2) it provides a correct $C(x^{\prime})$ but an incorrect $x^{\prime}$. In either case, the probability that it gets lucky and/or passes the challenge by cheating is $1/4$; after $r$ rounds, it is $(1/4)^{r}$. ∎
Theorem 5.2 (No-Data Algorithm Correctness Bound).
@@
(Sketch; a complete proof is in Appendix A.4). The accuracy of any evaluator is directly tied to its success in the EV protocol: if it knows the label, then the checks will pass and $\tilde{y}^{\prime}=y$. From Lemma 5.1, this holds with probability $1-(1/4)^{r}$. An application of the union bound yields the desired result.∎
Theorem 5.2 implies that a good evaluator with relatively high accuracy and success rate, and a well-calibrated $\phi$, will have a true accuracy higher than what the No-Data Algorithm outputs. Hence, reliable evaluators, as per the algorithm, may be used with the expectation that their ‘true’ performance will be better. Another interpretation of this result is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier.
Remark that the assumption for Theorem 5.2 for the evaluator always lying when mislabelling (and hence always being wrong) is not necessarily applicable to all evaluators. A particularly tricky evaluator could still guess the right label, yet not pass the challenges. We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust, especially in scenarios without any references.